# INACIA: Integrating Large Language Models in Brazilian Audit Courts: Opportunities and Challenges

### 19 Jan 2024 v2

@@ 5. Results
Table 2 presents the results of evaluating the recommendations generated by the INACIA pipeline using the checklist approach. This method assesses the alignment of the generated recommendations with the checklist of elements derived from the reference text. The table shows the mean, standard deviation (Std), minimum (Min), and maximum (Max) values for precision, recall, and the F1 score. These metrics provide insights into the consistency and completeness of the generated recommendations compared to the reference text. The average precision score of 0.592 indicates that, on average, 59.2% of the checklist elements were found in the generated recommendations. This suggests a moderate level of accuracy in capturing the necessary elements in the recommendations. The average recall score is 0.401, indicating that roughly 40.1% of the reference text elements were present in the generated recommendations. This lower recall suggests that the generated recommendations do not capture significant reference text elements. The average F1 score, the harmonic mean of precision and recall, is 0.429. This score reflects a balance between precision and recall but suggests room for improvement in both aspects. The standard deviation values for precision (0.212), recall (0.275), and F1 (0.225) show a certain degree of variability in the performance across different cases. Overall, these results suggest that while there are instances where the recommendations closely match the reference text, there is a noticeable variation in the quality of the generated recommendations. The relatively lower recall scores indicate a need for improvement in capturing all the necessary elements from the reference text in the generated recommendations.
Although the INACIA pipeline does not always produce the expected results, it is possible to follow the reasoning of the INACIA pipeline to understand the decisions made by the agent. This helps the auditor to find the evidence. For example, the INACIA pipeline may not classify the case as admissible regarding some criterion in the admissibility examination. However, it is possible to follow the reasoning of the INACIA pipeline to understand why the case was not classified as admissible. This helps the auditor to find the evidence that supports the decision made by INACIA.
@@ 6. Conclusions
In this paper, we presented INACIA, an innovative system designed to integrate Large Language Models (LLMs) into the Brazilian Federal Court of Accounts operational framework. Our experiments and analysis reveal that INACIA holds significant potential to transform and optimize the audit process by automating and streamlining case analysis and decision-making.
@@
Moving forward, we will focus on enhancing INACIA’s capabilities through continuous learning and adaptation. Besides, we will explore implementing an auditing system based on INACIA’s architecture, allowing us to evaluate the system’s performance in a real-world setting. In conclusion, INACIA represents a significant step forward in applying artificial intelligence in the legal domain. Its development and implementation in the Brazilian audit courts can serve as a blueprint for similar initiatives globally.

### 26 Feb 2024 v3

@@ 6. Results
Table 2 presents the results of evaluating the recommendations generated by the INACIA pipeline using the checklist approach. This method assesses the alignment of the generated recommendations with the checklist of elements derived from the reference text. The table shows the mean, standard deviation (Std), minimum (Min), and maximum (Max) values for precision, recall, and the F1 score. These metrics provide insights into the consistency and completeness of the generated recommendations compared to the reference text. The average precision score of 0.592 indicates that, on average, 59.2% of the checklist elements were found in the generated recommendations. This suggests a moderate level of accuracy in capturing the necessary elements in the recommendations. The average recall score is 0.401, indicating that roughly 40.1% of the reference text elements were present in the generated recommendations. This lower recall suggests that the generated recommendations do not capture significant reference text elements. The average F1 score is 0.429 which reflects a balance between precision and recall but suggests room for improvement in both aspects. The standard deviation values for precision (0.212), recall (0.275), and F1 (0.225) show a certain degree of variability in the performance across different cases. Overall, these results suggest that while there are instances where the recommendations closely match the reference text, there is a noticeable variation in the quality of the generated recommendations. The relatively lower recall scores indicate a need for improvement in capturing all the necessary elements from the reference text in the generated recommendations.
@@
Although the INACIA pipeline does not always produce the expected results, it is possible to follow the reasoning of the INACIA pipeline to understand the decisions made by the agent. This helps the auditor to find the evidence. For example, the INACIA pipeline may not classify the case as admissible regarding some criterion in the admissibility examination. However, it is possible to follow the reasoning of the INACIA pipeline to understand why the case was not classified as admissible. Figure 8 shows an example of the reasoning produced by INACIA for the admissibility examination considering the public interest criterion. The reasoning explains why the case was classified as admissible regarding the public interest criterion. This helps the auditor to understand the decision made by INACIA and find the evidence that supports it.
Explicação: As alegações do autor da representação indicam a existência de possíveis irregularidades no Pregão Eletrônico nº 03/2021 promovido pelo IBAMA. Entre as irregularidades apontadas, estão a não observância das exigências do instrumento convocatório e das leis vigentes da licitação pública, atos praticados por servidores do órgão que podem ser considerados obscuros e que ferem o caráter competitivo do certame, e a habilitação de uma empresa que não apresentou a certidão de falência exigida pelo edital.
Caso essas alegações sejam confirmadas, elas podem indicar a existência de interesse público, conforme definido no Manual de Representações e Denúncias do TCU. Isso porque, se comprovadas, essas irregularidades podem representar um potencial risco de dano ao erário, configurar uma restrição indevida à competição, representar um potencial risco de inexecução ou execução insatisfatória do contrato, configurar uma fraude à licitação ou ao contrato, e representar um potencial risco de ineficácia na execução de política pública.
Admissível: Sim
(a) Brazilian Portuguese version.
Explanation: The allegations made by the author of the representation indicate the existence of possible irregularities in Electronic Auction No. 03/2021 promoted by IBAMA. Among the irregularities highlighted are non-compliance with the requirements of the call for proposals and the current laws governing public bidding, acts carried out by the agency’s employees that may be considered obscure and that harm the competitive nature of the event, and the qualification of a company that does not present the bankruptcy certificate required by the notice.
If these allegations are confirmed, they may indicate the existence of public interest, as defined in the TCU Representations and Complaints Manual. This is because, if proven, these irregularities may represent a potential risk of damage to the treasury, constitute an undue restriction on competition, represent a potential risk of non-execution or unsatisfactory execution of the contract, constitute fraud in the bid or the contract, and represent a potential risk of ineffectiveness in the execution of public policy.
Admissible: Yes
(b) English version.
Figure 8. Example of the reasoning produced by INACIA for the admissibility examination considering the public interest criterion.
7. Discussion
7.1. Limitations
INACIA departs from common, simple uses of large language models in a few ways: (1) it uses long context windows (and retrieval-augmented generation) techniques to extract information spread across multiple case documents; (2) it requires the LLM to reason through many, often conflicting documents and identify the most important claims being made and present those to its user (e.g., evidence presented by plaintiff vs. jurisprudence); (3) it allows the LLM to search for external documents that are relevant to each analysis before generating text for the user; (4) it produces a list of adjudication recommendations with little knowledge of what information takes precedence (e.g., is it the law or jurisprudence, is it a more recent or less recent piece of legislation, is it one claim or all claims, etc.); (5) its main product is free-form text with recommendations that are not easily compared to purported ground truth retrieved from work done by TCU auditors. Naturally, thus, the system has limitations due to the inability of LLMs to perform all the tasks above. Besides these inherent limitations, there are practical limitations akin to large projects such as INACIA; e.g., the time constraints on developing each module and then integrating the whole application; the balance between developing a very advanced, cutting-edge but also unstable technology; the gap in business, context-specific knowledge between the research team and TCU auditors, etc. Despite these limitations, we firmly believe INACIA has great potential to assist TCU staff in their tasks and, therefore, add a lot of value to the TCU case processing structure.
7.2. Ethical Implications
Whenever artificial intelligence systems are being developed in real-world applications, one of the first concerns of decision-makers responsible for the system’s adoption is whether the models are biased or fair. While acknowledging these as real issues in any AI system, we believe INACIA to be relatively shielded against such concerns due to the nature of the case processing steps and information provided to the language models to complete each step in the INACIA pipeline. First and foremost, the complaints INACIA processes are not filed against individual defendants nor any individual trait is used in the analysis of case components. Secondly, the techniques used to prompt INACIA are all grounding its answers onto information belonging to TCU’s knowledge base, such as jurisprudence, legal statutes and norms, and previous cases handled by the audit court. This structure ensures that INACIA’s adjudication recommendations are based on established legal principles and precedents, significantly mitigating the risks of bias.
Another important ethical consideration is the transparency, explainability, and accountability behind INACIA’s decision-making process. Since the potential consequences of INACIA’s recommendations can be large, it is imperative that the reasoning behind these decisions is clear and understandable. We believe this risk to be limited since none of the decision factors used in INACIA prompts contains elements departing from current adjudication practices at TCU. TCU staff has provided extensive support and auditing materials so that INACIA could replicate as close as possible the work of a TCU auditor. Secondly, TCU auditors and eventually the court’s judicial panel bear ultimate responsibility for any case decision. TCU has not removed, nor plans on removing, any human oversight of INACIA’s outputs.999 There are other relatively minor risks that any new system carries, such as data security and privacy and job displacement, that are not particular to INACIA. As far as data security and privacy go, INACIA is being built to be compatible with existing TCU systems without adding any new risk. As to potential job displacement risks, TCU is precisely looking to replace tasks in its staff’s work assignments with AI such that staff can shift their skills to other, more important tasks within TCU.
@@ 8. Conclusions
In this paper, we presented INACIA, an innovative system designed to integrate Large Language Models (LLMs) into the Brazilian Federal Court of Accounts operational framework. Our experiments and analyses reveal that INACIA holds significant potential to support the auditor in the audit process with greater speed and security in analyzing cases and making decisions.
@@
Moving forward, we will focus on enhancing INACIA’s capabilities through continuous development and improvements. Besides, we will explore implementing an auditing system based on INACIA’s architecture, allowing us to evaluate the system’s performance in a real-world setting. In conclusion, INACIA represents a significant step forward in applying artificial intelligence in the legal domain. Its development and implementation in the Brazilian audit courts can serve as a blueprint for similar initiatives globally.
Appendix A List of Variables in the Validation Dataset
Table 3. List of variables in the validation dataset.
Variable Name Description
Basic Information Extraction
numero_processo
Individual case ID
info_uasg
Federal agency under investigation (unique ID)
info_author_name
Plaintiffś name
info_author_doc
Plaintiffś document ID
info_author_doc_type
Plaintiffś document type
info_author_doc_complement_type
Plaintiff’s additional document type information
info_author_doc_complement
Plaintiff’s additional document information
info_case_summary
Case summary
info_case_jurisdictional_unit
Federal agency under investigation (name)
info_case_subject
Case subject
info_case_dispute_validity
Period during which procurement call remained open for proposals
info_case_type
TCU case classification
info_case_dispute_mode
Procurement awarding mechanism
info_case_dispute_criterion
Procurement awarding mechanism criteria
info_case_dispute_value
Procurement contract value
info_case_dispute_value_type
Procurement contract value estimated or observed
info_case_status
Procurement status
info_case_status_ambiguous
Procurement status ambiguous
info_procurement_id
Procurement ID
info_procurement_phase
Procurement stage
info_procurement_type
Procurement type
info_procurement_type_id
Procurement type ID
info_procurement_law
Law applicable to procurement call
Periculum in mora
cautelar_perigo_demora
Periculum in mora present (per TCU auditor evaluation)
cautelar_perigo_demora_rev
Reverse periculum in mora present (per TCU auditor evaluation)
Fumus boni iuris
cautelar_plausibilidade
Fumus boni iuris (per TCU auditor evaluation)
Admissibility Examination
admiss_legitimidade
Legitimacy confirmation
admiss_linguagem
Clear writing and language confirmation
admiss_indicio
Existence of evidence confirmation
admiss_competencia
Competency confirmation
admiss_interesse_pub
Existence of public interest
Allegations and Requests
alegacoes_txt
List of allegations
Recommendations
encaminhamentos_txt
List of recommendations made by TCU auditor